OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(1/2, 1) and T(1/2,4) X = ?

OpenStudy (anonymous):

find the slope of this line using \[y_{2} - y _{1}\over x_{2}- x _{1}\]then plug in the values for the point slope formula \[y-y _{1}=m(x-x _{1})\]where: y1= first y-value x1= first x-value m = slope

OpenStudy (anonymous):

once youve found the slope and worked through the point slope formula turn it into standard form :)

OpenStudy (anonymous):

That's the thing...when I try to use the slope formula, I'm unable to continue on, due to the simple fact I can't divide by 0

OpenStudy (anonymous):

then the slope of the line is undefined :) thus x=0

OpenStudy (anonymous):

i think you wrote the problem wrong or something....because this doesnt seem right

OpenStudy (anonymous):

No, this is exactly how it's written.

OpenStudy (anonymous):

then im not sure :\

OpenStudy (anonymous):

Oh wow! It was 1/2 XD

OpenStudy (anonymous):

what was? :)

OpenStudy (anonymous):

The slope. Using the second formula, I got y - 1/2 = m(x - 1/2) So I figured out it was 1/2

OpenStudy (anonymous):

1/2 is the x value for both points...

OpenStudy (anonymous):

Yeah, geometry likes to play by its own rules, huh? :P Well, that's apparently the correct answer, so...thanks for the help!

OpenStudy (anonymous):

umm wow okay :D :P ...cool well glad i helped :)

OpenStudy (anonymous):

msg if u have any other questions :)